OpenStudy (anonymous):

Line L has equation 2x - 3y = 5. Line M passes through the point (3, -10) and is parallel to line L. Determine the equation for line M.

OpenStudy (amistre64):

same 'equation' ; different sum...

OpenStudy (amistre64):

2x -3y = ?? ; when (x=3, y=-10)

OpenStudy (amistre64):

2(3) -3(-10) = ?? 6 +30 = 36 2x -3y = 36

OpenStudy (anonymous):

Thanks!:]

OpenStudy (amistre64):

youre welcome :)